OpenStudy (anonymous):

use rational exponents to write x^1/6*y^1/3*z^1/5

jhonyy9 (jhonyy9):

x*1 y*1 z*1 xyz --- x --- x --- = (x/6)(y/3)(z/5) = ---- 6 3 5 90